Hi I need help on this question I'm struggling with. Thank you in advance!
The question is: The area of a parking lot is 2475 ft.². The rectangular parking lot has dimensions such that the length is 10 feet longer than the width. What are the dimensions of the parking lot?

Answer by FrankM(1040)

let X be the average of the 2. So X-5 is the length, X+5, the width.
(X+5)(X-5) = 2475
X^2-25 = 2475
X^2 = 2500
X = 50
The numbers are 45 and 55.